# Learning Representations for Independence Testing

### v2

@@ B.5 Main Results
Employ Propositions B.8 and B.10 as uniform bounds on $|\hat{T}_{1}-T_{1}|$ and $|\hat{\sigma}^{2}-\sigma^{2}|$ respectively, then take the union bound to give us, with probability at least $1-\delta$,
@@
Term (b) can be handled in a similar manner, except we use Proposition B.9 as a uniform bound on $|\hat{T}_{0}-T_{0}|$. It turns out this gives us the same asymptotic bound as (a).
@@
Propositions B.10 and B.11 are uniform bounds on $|\hat{\sigma}^{2}-\sigma^{2}|$ and $|\hat{\xi}^{2}-\xi^{2}|$. Then, taking the union bound gives us, with probability at least $1-\delta$,
@@
Theorem B.13 (Consistency [van der Vaart, 1998, Theorem 5.7]).
@@
[…]
Appendix C Uniform Convergence: HSIC
C.1 Preliminaries
@@ C.2 Assumptions
C.3 Lemmas
Lemma C.1 (Concentration of $\hat{\sigma}_{\omega,\gamma}^{2}$).
For any kernels $k$ and $l$ satisfying Assumption (A’), with probability at least $1-\delta$ we have
@@ C.3 Lemmas
We apply McDiarmid’s inequality to $\hat{\sigma}^{2}$. First, we show that the variance estimator satisfies bounded differences. For convenience, we denote $(i,j,q,r)\in{\bf i}^{n}_{4}$ simply as $(i,j,q,r)$, and $(i,j,q)\backslash k$ to be the set of 3-tuples drawn without replacement from ${\bf i}^{n}_{3}$ that exclude the number $k$. Recall that
$\hat{\sigma}^{2}=16\left(\frac{1}{(n)_{4}(n-1)_{3}}\sum_{\begin{subarray}{c}(i,j,q,r)\\ (b,c,d)\backslash i\end{subarray}}H_{ijqr}H_{ibcd}-\hat{\eta}^{2}\right).$
@@
We can bound the second term using $|\hat{\eta}|\leq 4\nu_{k}\nu_{l}$ and the bounded difference result (15) from Proposition C.3:
@@
Lemma C.2 (Bias of $\hat{\sigma}_{\omega,\gamma}^{2}$).
For any kernels $k$ and $l$ satisfying Assumption (A’), the bias is bounded by
@@
$\operatorname*{\mathbb{E}}\hat{\sigma}^{2}=16\left(\frac{1}{(n)_{4}(n-1)_{3}}\sum_{\begin{subarray}{c}(i,j,q,r)\\ (b,c,d)\backslash i\end{subarray}}\operatorname*{\mathbb{E}}[H_{ijqr}H_{ibcd}]-\frac{1}{(n)^{2}_{4}}\sum_{\begin{subarray}{c}(i,j,q,r)\\ (a,b,c,d)\end{subarray}}\operatorname*{\mathbb{E}}[H_{ijqr}H_{abcd}]\right).$
@@
$|\operatorname*{\mathbb{E}}\hat{\sigma}^{2}-\sigma^{2}|$ $=16\left|\frac{1}{(n)_{4}(n-1)_{3}}\sum_{\begin{subarray}{c}(i,j,q,r)\\ (b,c,d)\backslash i\end{subarray}}\operatorname*{\mathbb{E}}[H_{ijqr}H_{ibcd}]-\operatorname*{\mathbb{E}}[H_{1234}H_{1567}]-\frac{1}{(n)^{2}_{4}}\sum_{\begin{subarray}{c}(i,j,q,r)\\ (a,b,c,d)\end{subarray}}\operatorname*{\mathbb{E}}[H_{ijqr}H_{abcd}]+\operatorname*{\mathbb{E}}[H_{1234}H_{5678}]\right|$
@@
C.4 Propositions
Proposition C.3 (Uniform convergence of $\widehat{\operatorname{HSIC}}_{u}$).
Under Assumptions (A’), (B’) and (C’), we have that with probability at least $1-\delta$,
@@ C.4 Propositions
$|H^{(\theta)}_{1234}-H^{(\theta^{\prime})}_{1234}|$ $\leq\frac{1}{4!}\sum^{(1234)}_{(abcd)}\left|K^{(\omega)}_{ab}(L^{(\gamma)}_{ab}+L^{(\gamma)}_{cd}-2L^{(\gamma)}_{ac})-K^{(\omega^{\prime})}_{ab}(L^{(\gamma^{\prime})}_{ab}+L^{(\gamma^{\prime})}_{cd}-2L^{(\gamma^{\prime})}_{ac})\right|$
$\leq\frac{1}{4!}\sum^{(1234)}_{(abcd)}\left(\left|K^{(\omega)}_{ab}L^{(\gamma)}_{ab}-K^{(\omega^{\prime})}_{ab}L^{(\gamma^{\prime})}_{ab}\right|+\left|K^{(\omega)}_{ab}L^{(\gamma)}_{cd}-K^{(\omega^{\prime})}_{ab}L^{(\gamma^{\prime})}_{cd}\right|+2\left|K^{(\omega^{\prime})}_{ab}L^{(\gamma^{\prime})}_{ac}-K^{(\omega)}_{ab}L^{(\gamma)}_{ac}\right|\right).$
@@
$\left|K^{(\omega)}_{ab}L^{(\gamma)}_{ab}-K^{(\omega^{\prime})}_{ab}L^{(\gamma^{\prime})}_{ab}\right|$ $=\left|K^{(\omega)}_{ab}L^{(\gamma)}_{ab}-K^{(\omega)}_{ab}L^{(\gamma^{\prime})}_{ab}+K^{(\omega)}_{ab}L^{(\gamma^{\prime})}_{ab}-K^{(\omega^{\prime})}_{ab}L^{(\gamma^{\prime})}_{ab}\right|$
$\leq\left|K^{(\omega)}_{ab}\right|\left|L^{(\gamma)}_{ab}-L^{(\gamma^{\prime})}_{ab}\right|+\left|L^{(\gamma^{\prime})}_{ab}\right|\left|K^{(\omega)}_{ab}-K^{(\omega^{\prime})}_{ab}\right|$
@@
This expression is true for all three components of $|H^{(\theta)}_{1234}-H^{(\theta^{\prime})}_{1234}|$ and so it follows that
@@
Now, we study the random error function $\Delta:=\hat{\eta}-\eta$. Note that $\operatorname*{\mathbb{E}}\Delta=0$ since $\hat{\eta}$ is unbiased, and $|H_{ijqr}|\leq 4\nu_{k}\nu_{l}$ via Assumption (A’). This $\hat{\eta}$, and hence $\Delta$, satisfies bounded differences. Let $F$ denote the kernel tensor $H$ but with sample $(X_{\ell},Y_{\ell})$ replaced by $(X^{\prime}_{\ell},Y^{\prime}_{\ell})$ so that $F$ agrees with $H$ except at indicies $\ell$, and let $\hat{\eta}^{\prime}=\frac{1}{(n)_{4}}\sum_{(i,j,q,r)\in{\bf i}^{n}_{4}}F_{ijqr}$ be it’s HSIC estimator.
For convenience, we denote $(i,j,q,r)\in{\bf i}^{n}_{4}$ simply as $(i,j,q,r)$, and $(i,j,q)\backslash k$ to be the set of 3-tuples drawn without replacement from ${\bf i}^{n}_{3}$ that exclude the number $k$. We can compute the maximal bounded difference $|\Delta-\Delta^{\prime}|=|\hat{\eta}-\hat{\eta}^{\prime}|$ as
@@
$=\frac{1}{(n)_{4}}\left(\sum_{(j,q,r)\backslash\ell}\underbrace{|H_{\ell jqr}-F_{\ell jqr}|}_{\leq 8\nu_{k}\nu_{l}}+\sum_{(i,q,r)\backslash\ell}|H_{i\ell qr}-F_{i\ell qr}|+\sum_{(i,j,r)\backslash\ell}|H_{ij\ell r}-F_{ij\ell r}|+\sum_{(i,j,q)\backslash\ell}|H_{ijq\ell}-F_{ijq\ell}|\right)$
@@
Proposition C.4 (Uniform convergence of $\hat{\sigma}_{\omega,\gamma}^{2}$).
Under Assumptions (A’), (B’) and (C’), we have that with probability at least $1-\delta$,
@@
We use an $\epsilon$-net argument on both spaces $\Omega$ and $\Gamma$. Using the same construction as in Proposition C.3, we once again decompose our convergence bound:
@@
$|\sigma^{2}_{\theta}-\sigma^{2}_{\theta^{\prime}}|$ $=16\left|\operatorname*{\mathbb{E}}[H^{(\theta)}_{1234}H^{(\theta)}_{1567}-H^{(\theta^{\prime})}_{1234}H^{(\theta^{\prime})}_{1567}]-\operatorname*{\mathbb{E}}[H^{(\theta)}_{1234}H^{(\theta)}_{5678}]+\operatorname*{\mathbb{E}}[H^{(\theta^{\prime})}_{1234}H^{(\theta^{\prime})}_{5678}]]\right|$
$\leq 16\operatorname*{\mathbb{E}}\left|H^{(\theta)}_{1234}H^{(\theta)}_{1567}-H^{(\theta^{\prime})}_{1234}H^{(\theta^{\prime})}_{1567}\right|+16\operatorname*{\mathbb{E}}\left|H^{(\theta)}_{1234}H^{(\theta)}_{5678}-H^{(\theta^{\prime})}_{1234}H^{(\theta^{\prime})}_{5678}\right|.$
Under Assumptions (A’) and (C’) we know that $|H_{1234}|\leq 4\nu_{k}\nu_{l}$ and $|H^{(\theta)}_{1234}-H^{(\theta^{\prime})}_{1234}|\leq 4\nu_{k}L_{l}\rho_{\Gamma}+4\nu_{l}L_{k}\rho_{\Omega}$, and so
$|H^{(\theta)}_{1234}H^{(\theta)}_{1567}-H^{(\theta^{\prime})}_{1234}H^{(\theta^{\prime})}_{1567}|$ $\leq|H^{(\theta)}_{1234}H^{(\theta)}_{1567}-H^{(\theta)}_{1234}H^{(\theta^{\prime})}_{1567}|+|H^{(\theta)}_{1234}H^{(\theta^{\prime})}_{1567}-H^{(\theta^{\prime})}_{1234}H^{(\theta^{\prime})}_{1567}|$
$=|H^{(\theta)}_{1234}||H^{(\theta)}_{1567}-H^{(\theta^{\prime})}_{1567}|+|H^{(\theta^{\prime})}_{1567}||H^{(\theta)}_{1234}-H^{(\theta^{\prime})}_{1234}|$
@@
Next, using Lemma C.1 and Lemma C.2 followed by a union bound over the $T_{\Omega}T_{\Gamma}$ center combinations gives us, with probability at least $1-\delta$,
@@
Finally, we combine Equations 16, 17 and 18 to get
@@
C.5 Main Results
Theorem C.5 (Uniform convergence of $\hat{J}_{\text{HSIC}}^{\lambda}$).
Under Assumptions (A’), (B’) and (C’), let $\Theta\subseteq\Omega\times\Gamma$ be the set of kernel parameters $\theta\in\Theta$ for which $\sigma_{\theta}^{2}\geq s^{2}$, and take $\lambda=n^{-1/3}$. Assume $\nu_{k},\nu_{l}\geq 1$. Then, with probability at least $1-\delta$,
@@ C.5 Main Results
Proposition C.3 and Proposition C.4 show the uniform convergence of $\hat{\eta}_{\theta}$ and $\hat{\sigma}_{\theta}$, from which we get that with probability at least $1-\delta$, the error is at most
@@
Appendix D Experiment Details
@@
D.1 Training & Test Details
@@
Regarding the power vs. test size experiments, we use a training size of 10,000 for HDGM $\leq$ 30, 100,000 for HDGM $>$ 30, and 2000 validation samples for all dimensions. For the Sinusoid problem, we train on 5,000 samples and use 1,000 for validation. RatInABox uses a training size of 4,000 samples without validation, and Wine uses 1,200 training samples without validation.
@@
D.2 Architectures
@@
High-dimensional Gaussian mixture. We use a feed-forward network for our deep kernel featurizer with latent dimensions $2d,3d,\text{and }2d$. Details of each model is given in Table 1.
@@
RatInABox. We use a feed-forward featurizer with details given in Table 2. Unlike the previous two problems, the sample spaces $\mathcal{X}$ and $\mathcal{Y}$ are not equivalent, and so the deep featurizers for $k$ and $l$ have different architectures.
@@
Appendix E Additional Experiments
E.1 High-Dimensional Gaussian Mixture
We provide comprehensive power versus test size results for the HDGM problem at all dimensions $d=\{2,4,5,10,15,20\}$ in Figure 9. Overall, our method HSIC-D/Dx achieves highest power at the smallest test sizes.
@@ E.1 High-Dimensional Gaussian Mixture
These are baselines considered in this paper. See more details in Section 6.
@@
Additionally, we demonstrate the effectiveness of our method at various dimensions $d/2$ by examining the empirical test power at HDGM-$d$ for $d\in\{4,8,10,20,30,40,50\}$ with fixed test sizes $m$. Results are shown in Figure 10. Again, HSIC-D exhibits the highest test power across all dimensions. When using a small number of test samples (e.g. $m=100$), the performance of HSIC-D slightly degrades with increasing dimension, whereas at larger test sample sizes it consistently has near-perfect power.
@@
These are baselines considered in this paper. See more details in Section 6.
@@
E.2 Type-I Error
Table 3 shows that the type-I error rates for our optimization-based tests are well-controlled.
@@ E.2 Type-I Error
E.3 Maximizing HSIC vs. its SNR
We examine the trade-off between directly optimizing HSIC versus its SNR objective $J$. The results on power versus test size are shown in Figure 11. Optimizing our proposed objective $J$ significantly outperforms optimizing HSIC for all problems.
@@
E.4 Independence Testing with MMD
Let =$\mathbb{Z}=\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ be a test set and $\mathcal{S}_{m}$ be the permutation group of $[m]$ with elements $\sigma\in\mathcal{S}_{m}$. Suppose we take $p$ permutations $\boldsymbol{\sigma}=\{\sigma_{1},...,\sigma_{p}\}$, with each permutation sampled uniformly from $\mathcal{G}_{m}$. We define the action of ${\boldsymbol{\sigma}}$ on test samples $\mathbb{Z}$ as ${\boldsymbol{\sigma}}\mathbb{Z}=\{(X_{i},Y_{\sigma_{\ell}(i)})\}_{i\in[m],\ell\in[p]}$, and the action of ${\boldsymbol{\sigma}}$ on the empirical distribution ^$\hat{\mathbb{P}}_{XY}$ as the empirical distribution of the permuted samples, i.e. ${\boldsymbol{\sigma}}\hat{\mathbb{P}}_{XY}=\frac{1}{np}\sum_{i=1}^{n}\sum_{\ell=1}^{p}\delta_{X_{i}\times Y_{\sigma_{\ell}(i)}}$.
One way we can construct an independence criterion is by taking the MMD between $\mathbb{P}_{XY}$ and $\mathbb{P}_{X}\times\mathbb{P}_{Y}$. We do this for MMD-D by using the empirical distributions ^$\hat{\mathbb{P}}_{XY}$ and ${\boldsymbol{\sigma}}\hat{\mathbb{P}}_{XY}$ respectively, which yields

### v3

@@ C.5 Main Results
Employ Propositions˜C.8 and C.10 as uniform bounds on $|\hat{T}_{1}-T_{1}|$ and $|\hat{\sigma}^{2}-\sigma^{2}|$ respectively, then take the union bound to give us, with probability at least $1-\delta$,
@@
Term (b) can be handled in a similar manner, except we use Proposition˜C.9 as a uniform bound on $|\hat{T}_{0}-T_{0}|$. It turns out this gives us the same asymptotic bound as (a).
@@
Propositions˜C.10 and C.11 are uniform bounds on $|\hat{\sigma}^{2}-\sigma^{2}|$ and $|\hat{\xi}^{2}-\xi^{2}|$. Then, taking the union bound gives us, with probability at least $1-\delta$,
@@
Theorem C.13 (Consistency [van der Vaart, 1998, Theorem 5.7]).
@@
is a consistent estimator, i.e. $\hat{\theta}_{n}\xrightarrow[]{p}\theta^{*}$.
Appendix D Uniform Convergence: HSIC
D.1 Preliminaries
@@ D.2 Assumptions
D.3 Lemmas
Lemma D.1 (Concentration of $\hat{\sigma}_{\omega,\gamma}^{2}$).
For any kernels $k$ and $l$ satisfying Assumption˜(A’), with probability at least $1-\delta$ we have
@@ D.3 Lemmas
We apply McDiarmid’s inequality to $\hat{\sigma}^{2}$. First, we show that the variance estimator satisfies bounded differences. For convenience, we denote $(i,j,q,r)\in{\bf i}^{n}_{4}$ simply as $(i,j,q,r)$, and $\k(i,j,q)\backslash$ k to be the set of 3-tuples drawn without replacement from ${\bf i}^{n}_{3}$ that exclude the number $k$. Recall that
[…]
@@
We can bound the second term using $|\hat{\eta}|\leq 4\nu_{k}\nu_{l}$ and the bounded difference result (15) from Proposition˜D.3:
@@
$|\hat{\sigma}^{2}-\operatorname*{\mathbb{E}}\hat{\sigma}^{2}|\leq 6144\nu^{2}_{k}\nu^{2}_{l}\sqrt{\frac{2}{n}\log\frac{2}{\delta}}.\qed$
Lemma D.2 (Bias of $\hat{\sigma}_{\omega,\gamma}^{2}$).
For any kernels $k$ and $l$ satisfying Assumption˜(A’), the bias is bounded by
@@
[…]
@@
[…]
@@
D.4 Propositions
Proposition D.3 (Uniform convergence of $\widehat{\operatorname{HSIC}}_{u}$).
Under Assumptions˜(A’), (B’) and (C’), we have that with probability at least $1-\delta$,
@@ D.4 Propositions
[…]
@@
[…]
@@
[…]
@@
Now, we study the random error function $\Delta:=\hat{\eta}-\eta$. Note that $\operatorname*{\mathbb{E}}\Delta=0$ since $\hat{\eta}$ is unbiased, and $|H_{ijqr}|\leq 4\nu_{k}\nu_{l}$ via Assumption˜(A’). This $\hat{\eta}$, and hence $\Delta$, satisfies bounded differences. Let $F$ denote the kernel tensor $H$ but with sample $(X_{\ell},Y_{\ell})$ replaced by $(X^{\prime}_{\ell},Y^{\prime}_{\ell})$ so that $F$ agrees with $H$ except at indicies $\ell$, and let $\hat{\eta}^{\prime}=\frac{1}{(n)_{4}}\sum_{(i,j,q,r)\in{\bf i}^{n}_{4}}F_{ijqr}$ be it’s HSIC estimator.
For convenience, we denote $(i,j,q,r)\in{\bf i}^{n}_{4}$ simply as $(i,j,q,r)$, and $\k(i,j,q)\backslash$ k to be the set of 3-tuples drawn without replacement from ${\bf i}^{n}_{3}$ that exclude the number $k$. We can compute the maximal bounded difference $|\Delta-\Delta^{\prime}|=|\hat{\eta}-\hat{\eta}^{\prime}|$ as
@@
[…]
@@
Proposition D.4 (Uniform convergence of $\hat{\sigma}_{\omega,\gamma}^{2}$).
Under Assumptions˜(A’), (B’) and (C’), we have that with probability at least $1-\delta$,
@@
We use an $\epsilon$-net argument on both spaces $\Omega$ and $\Gamma$. Using the same construction as in Proposition˜D.3, we once again decompose our convergence bound:
@@
[…]
Under Assumptions˜(A’) and (C’) we know that $|H_{1234}|\leq 4\nu_{k}\nu_{l}$ and $|H^{(\theta)}_{1234}-H^{(\theta^{\prime})}_{1234}|\leq 4\nu_{k}L_{l}\rho_{\Gamma}+4\nu_{l}L_{k}\rho_{\Omega}$, and so
[…]
@@
Next, using Lemma˜D.1 and Lemma˜D.2 followed by a union bound over the $T_{\Omega}T_{\Gamma}$ center combinations gives us, with probability at least $1-\delta$,
@@
Finally, we combine Equations˜16, 17 and 18 to get
@@
D.5 Main Results
Theorem D.5 (Uniform convergence of $\hat{J}_{\text{HSIC}}^{\lambda}$).
Under Assumptions˜(A’), (B’) and (C’), let $\Theta\subseteq\Omega\times\Gamma$ be the set of kernel parameters $\theta\in\Theta$ for which $\sigma_{\theta}^{2}\geq s^{2}$, and take $\lambda=n^{-1/3}$. Assume $\nu_{k},\nu_{l}\geq 1$. Then, with probability at least $1-\delta$,
@@ D.5 Main Results
Proposition˜D.3 and Proposition˜D.4 show the uniform convergence of $\hat{\eta}_{\theta}$ and $\hat{\sigma}_{\theta}$, from which we get that with probability at least $1-\delta$, the error is at most
@@
Appendix E Experimental Details
@@
E.1 Training & Test Details
@@
Regarding the power vs. test size experiments, we use a training size of 10,000 for HDGM $\leq$ 30, 100,000 for HDGM >> 30, and 2000 validation samples for all dimensions. For the Sinusoid problem, we train on 5,000 samples and use 1,000 for validation. RatInABox uses a training size of 4,000 samples without validation, and Wine uses 1,200 training samples without validation.
@@
E.2 Validation
Although validation sets were only used for early stopping in our experiments, it’s certainly possible to use them for hyperparameter selection. The validity of permutation tests still hold as long as this validation set is separate from the test set. As for the validation criterion, a natural approach is to evaluate the same approximation of the asymptotic power used for our training objective. This avoids the need for relatively-expensive permutations, while providing a reasonable estimate of performance. Alternatively, we could also directly estimate the test statistic in cases where there isn’t a simple expression for asymptotic power, though this approach may be less reliable since the statistic is not necessarily correlated with test power. It is also possible to select hyperparameters through cross-validation, where the training data is split into folds: one fold for evaluating hyperparameters and the remaining folds for training the model. Again, as long as this data is separate from the test set, any permutation test would be well-controlled. The drawback to these validation-based approaches is that effective power is reduced since we need to construct this validation set from either the training or test data.
Ideally, we would like to perform hyperparameter selection while avoiding sample splitting, but this is a challenging problem. One potential approach is to perform separate tests corresponding to each hyperparameter, and then combine these tests using a multiple test correction procedure like MMDAgg (Schrab et al., 2023) to ensure valid level control on the test data. However, this is potentially expensive if hyperparameters need to be optimized. It’s also unclear if this would perform better than simply using a validation set due to the conservatism of the multiple test correction procedure. A related approach is MMDFuse (Biggs et al., 2023), which constructs an aggregate statistic by combining MMD estimates at varying bandwidths. This approach could be naturally extended to consider estimates based on different training sets and would naturally fit into our optimization framework; though doing so in a way that does not explode compute nor reduce critic quality is another challenge.
E.3 Architectures
@@
High-dimensional Gaussian mixture. We use a feed-forward network for our deep kernel featurizer with latent dimensions $2d,3d,\text{and }2d$. Details of each model is given in Table˜1.
@@
RatInABox. We use a feed-forward featurizer with details given in Table˜2. Unlike the previous two problems, the sample spaces $\mathcal{X}$ and $\mathcal{Y}$ are not equivalent, and so the deep featurizers for $k$ and $l$ have different architectures.
@@
Appendix F Additional Experiments
F.1 High-Dimensional Gaussian Mixture
We provide comprehensive power versus test size results for the HDGM problem at all dimensions $d=\{2,4,5,10,15,20\}$ in Figure˜9. Overall, our method HSIC-D/Dx achieves highest power at the smallest test sizes.
@@ F.1 High-Dimensional Gaussian Mixture
[…]
@@
Additionally, we demonstrate the effectiveness of our method at various dimensions $d/2$ by examining the empirical test power at HDGM-$d$ for $d\in\{4,8,10,20,30,40,50\}$ with fixed test sizes $m$. Results are shown in Figure˜10. Again, HSIC-D exhibits the highest test power across all dimensions. When using a small number of test samples (e.g. $m=100$), the performance of HSIC-D slightly degrades with increasing dimension, whereas at larger test sample sizes it consistently has near-perfect power.
@@
[…]
@@
F.2 Type-I Error
Table˜3 shows that the type-I error rates for our optimization-based tests are well-controlled.
@@ F.2 Type-I Error
F.3 Maximizing HSIC vs. its SNR
We examine the trade-off between directly optimizing HSIC versus its SNR objective $J$. The results on power versus test size are shown in Figure˜11. Optimizing our proposed objective $J$ significantly outperforms optimizing HSIC for all problems.
@@
F.4 Independence Testing with MMD
Let $\mathbb{Z}=\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ be a test set and $\mathcal{S}_{m}$ be the permutation group of $[m]$ with elements $\sigma\in\mathcal{S}_{m}$. Suppose we take $p$ permutations $\boldsymbol{\sigma}=\{\sigma_{1},...,\sigma_{p}\}$, with each permutation sampled uniformly from $\mathcal{G}_{m}$. We define the action of $\boldsymbol{\sigma}$} on test samples $\mathbb{Z}$ as ${\boldsymbol{\sigma}}\mathbb{Z}=\{(X_{i},Y_{\sigma_{\ell}(i)})\}_{i\in[m],\ell\in[p]}$, and the action of $\boldsymbol{\sigma}$} on the empirical distribution $\hat{\mathbb{P}}_{XY}$ as the empirical distribution of the permuted samples, i.e. ${\boldsymbol{\sigma}}\hat{\mathbb{P}}_{XY}=\frac{1}{np}\sum_{i=1}^{n}\sum_{\ell=1}^{p}\delta_{X_{i}\times Y_{\sigma_{\ell}(i)}}$.
One way we can construct an independence criterion is by taking the MMD between $\mathbb{P}_{XY}$ and $\mathbb{P}_{X}\times\mathbb{P}_{Y}$. We do this for MMD-D by using the empirical distributions $\hat{\mathbb{P}}_{XY}$ and ${\boldsymbol{\sigma}}\hat{\mathbb{P}}_{XY}$ respectively, which yields